# Superpose Singular Features for Model Merging

### v1

@@ 3.2 Merge Task Matrices to Superpose Singular Features
where $\boldsymbol{\alpha}$ is the vector of variables $\alpha_{i^{\prime}}^{(k^{\prime})}$ and $\boldsymbol{\sigma}$ is the vector of singular values $\sigma_{i}^{(k)}$, with consistent indexing order over $i,k,i^{\prime},k^{\prime}$ across $\mathbf{\Sigma}$, $\mathbf{U}$ and $\mathbf{V}$. By solving the linear system in (4), we obtain the optimal weights $\boldsymbol{\alpha}^{*}$ for merging task matrices according to (1). The merged linear transformation matrix $\mathbf{P}$ is then obtained by adding the merged task matrix $\mathbf{M}$ to the pretrained matrix $\mathbf{P}_{\text{pre}}$ with a scaling factor $\gamma$: $\mathbf{P}=\mathbf{P}_{\text{pre}}+\gamma\mathbf{M}$. The algorithm for merging task matrices is summarized in Algorithm 1.
Algorithm 1 Merging task matrices: $\mathsf{{SFTM}}(\mathbf{M}_{i},i=1,\ldots,T)$
@@
  Prepare $\mathbf{\Sigma}$, $\mathbf{U}$, and $\mathbf{V}$;
  Solve $\mathbf{\Sigma}\circ\mathbf{U}\circ\mathbf{V}\boldsymbol{\alpha}=\boldsymbol{\sigma}$ to obtain $\boldsymbol{\alpha}^{*}$;
@@ 3.3 Complete Algorithm
Except the linear transformation matrices, there are other parameters in the model that need to be merged, such as biases, normalization parameters, embeddings, and convolutional layers. For biases and embeddings, we merge them with task arithmetic, i.e., adding them task vectors together. For normalization parameters, we merge them by averaging the normalization parameters of the individual tasks because the normalization can be seen as a linear transformation with diagonal matrix. For convolutional layers, we convert them into linear transformations by reshaping the convolutional kernels into a matrix and then merge them with the same method as the task matrices.
We also follow Ties-Merging (Yadav et al., 2024) and apply a trimming step. This involves keeping only the top $\eta$ parameters by magnitude while setting the remaining parameters to zero. This preprocessing step helps reduce noise and focus on the most significant parameters during merging. We present the complete algorithm for merging models in Algorithm 2.
Algorithm 2 Complete Algorithm for Model Merging
1:  Input: finetuned models $\{\theta_{1},\ldots,\theta_{T}\}$, pretrained model $\theta_{\text{pre}}$
@@
6:     $\mathbf{M}\leftarrow\mathsf{{SFTM}}(\mathbf{M}_{i},i=1,\ldots,T)$Algorithm 1}
@@ 3.4 Discussion
Time Complexity
Because there are $T$ tasks matrices of size $m\times n$, SVD of these matrix takes the time complexity $O(Tmn^{2})$. To merge $T$ task matrices, the linear system has $r$ equations and variables to solve, which takes the time complexity $O(r^{3})$. Therefore, the overall time complexity of SFTM for merging $T$ task matrices is $O(Tmn^{2}+r^{3})$. See Section 4.3 for the time consumption of SFTM for various models.
@@
In this paper, our focus is on merging task matrices instead of fine-tuning matrices. We have discovered that fine-tuned linear transformation matrices tend to have more shared features across tasks. Therefore, merging the fine-tuned linear transformation matrix is not as efficient as task matrix because certain overlapping direction of singular vectors correspond to these common features. On the contrary, the singular vectors of the task matrix contain features that are more specific to individual tasks, making it a more effective approach for merging. See the results of merging fine-tuned matrices in Section 4.3.
@@ 4.1 Experimental Setup
We evaluate SFTM across diverse fine-tuning methods, tasks, model architectures, and model size. For the fully fine-tuning, we evaluate vision tasks on ViT and NLP tasks on T5 following protocols of Ilharco et al. (2022a) and Yadav et al. (2024), respectively. For parameter-efficient fine-tuning (PEFT), we evaluate LoRA on GPT-2 for NLP tasks since it uses linear transformation matrices through low-rank adaptation, which aligns well with our approach, rather than vector parameters like IA3 (Liu et al., 2022) in previous research. To evaluate SFTM when model size scales up, we follow the setup of Du et al. (2024) to merge LLMs. Additionally, to evaluate the out-of-distribution robustness, we follow Yadav et al. (2024) to test it on NLP tasks.
@@
We compare SFTM against four established model merging approaches: (1) Averaging (Choshen et al., 2022; Wortsman et al., 2022), which computes the element-wise mean of individual models; (2) Task Arithmetic (Ilharco et al., 2022a), which merges by scaling and adding task vectors to the initial model; (3) Fisher Merging (Matena and Raffel, 2022), which approximates Fisher Information Matrix to weight parameters based on their importance for each task and combines the weighted parameters into the final merged model; (4) RegMean (Jin et al., 2022), which computes a closed-form solution to a least-squares regression problem that aims to minimize the distance between the merged model’s activations and the individual models’ activations. (5) Ties-Merging, which enhances merging by eliminating redundant parameters and resolving sign conflicts; and (6) PCB-Merging, which balances parameter competition through intra-task significance and inter-task similarity analysis. We also report results from individual fine-tuned models and a pre-trained model on all tasks.

### v2

@@ 3.3 Merging by Solving Linear System
By solving the linear system (3), we obtain the optimal weights $\boldsymbol{\alpha}^{*}$ for merging task matrices according to (2). The merged linear transformation matrix $\mathbf{P}$ is then obtained by adding the merged task matrix $\mathbf{M}$ to the pre-trained matrix $\mathbf{P}_{\text{pre}}$ with a scaling factor $\gamma$: $\mathbf{P}=\mathbf{P}_{\text{pre}}+\gamma\mathbf{M}$.
Algorithm 1 Merging task matrices: $\mathsf{{STF}}(\{\mathbf{M}_{i}\})$
@@
 Prepare $\mathbf{U}$, and $\mathbf{V}$;
 Solve $(\mathbf{U}\circ\mathbf{V})\boldsymbol{\alpha}=\boldsymbol{\sigma}$ to obtain $\boldsymbol{\alpha}^{*}$;
@@
The algorithm for merging task matrices is summarized in Algorithm˜1 and Fig.˜2. Because there are $T$ tasks matrices of size $m\times n$, SVD of these matrices takes the time complexity $O(Tmn^{2})$. To merge $T$ task matrices, the linear system has $r$ equations and variables to solve, which takes the time complexity $O(r^{3})$. Therefore, the overall time complexity of STF for merging $T$ task matrices is $O(Tmn^{2}+r^{3})$. See Section˜4.3 for the time consumption of model merging with STF for various models.
@@ 3.4 Complete Algorithm
Except the linear transformation matrices, there are parameters in the model that need to be merged, such as biases, normalization parameters, and embeddings. For biases and embeddings, we merge them with task arithmetic, i.e., adding their task vectors together. For normalization parameters, we merge them by averaging the normalization parameters of the individual tasks because the normalization can be seen as a linear transformation with diagonal matrix.
We also follow Ties-Merging Yadav et al. (2023) and apply a trimming step. This involves keeping only the top $\eta\%$ parameters by magnitude while setting the remaining parameters to zero. This preprocessing step helps reduce noise and focus on the most significant parameters during merging. We present the complete algorithm for merging models in Algorithm˜2.
Algorithm 2 Complete Merging Algorithm
1:Input: fine-tuned models $\{\theta_{1},\ldots,\theta_{T}\}$, pre-trained model $\theta_{\text{pre}}$
@@
6:  $\mathbf{M}\leftarrow\mathsf{{STF}}(\{\mathbf{M}_{i}\}_{i=1}^{T})$ {Algorithm 1}
@@ 3.5 Discussion
In this paper, our focus is on merging task matrices instead of fine-tuning matrices. We have discovered that fine-tuned linear transformation matrices tend to have more shared features across tasks. Therefore, merging the fine-tuned linear transformation matrix is not as efficient as merging task matrix because certain overlapping direction of singular vectors correspond to these common features. On the contrary, the singular vectors of the task matrix contain features that are more specific to individual tasks, making it a more effective approach for merging. See the results of merging fine-tuned matrices in Section˜4.3.
Feature Interference and Scaling Factor
In STF, we scale the merged matrix $\mathbf{M}$ to reduce interference between task-specific features. During merging, features from different tasks can interfere since we combine them through linear superposition. While using singular vectors as basis helps minimize interference within each task, some features may still be diminished when their directions conflict with features from other tasks. The scaling factor $\gamma$ plays a crucial role in managing this tradeoff - a small $\gamma$ reduces interference but may weaken task-specific features, while a large $\gamma$ better preserves individual task capabilities but risks amplifying interference between tasks. Finding the optimal $\gamma$ requires careful tuning based on the specific tasks and model architecture. In Section˜5, we discuss future directions for reducing feature interference beyond scaling.
@@ 4.1 Experimental Setup
We evaluate STF across diverse fine-tuning methods, tasks, model architectures, and model size. For the fully fine-tuning, we evaluate NLP tasks on T5 following protocols of Yadav et al. (2023). For parameter-efficient fine-tuning (PEFT), we evaluate LoRA on GPT-2 for NLP tasks since it uses linear transformation matrices through low-rank adaptation, which aligns well with our approach, rather than vector parameters like IA3 Liu et al. (2022) in previous research. To evaluate STF when model size scales up, we follow the setup of Du et al. (2024) to merge LLMs. Additionally, to evaluate the out-of-distribution robustness, we follow Yadav et al. (2023) to test it on NLP tasks.
All experiments are conducted on NVIDIA A100 GPUs and AMD EPYC 7763 CPUs. STF is implemented in PyTorch and performed on NVIDIA A100 GPUs. The hyperparameter settings for STF are in Section˜D.2.
@@
We compare STF against four established model merging approaches: (1) Averaging Choshen et al. (2022); Wortsman et al. (2022), which computes the element-wise mean of individual models; (2) Task Arithmetic Ilharco et al. (2022), which merges by scaling and adding task vectors to the initial model; (3) Fisher Merging Matena and Raffel (2022), which approximates Fisher Information Matrix to weight parameters based on their importance for each task and combines the weighted parameters into the final merged model; (4) RegMean Jin et al. (2022), which computes a closed-form solution to a least-squares regression problem that aims to minimize the distance between the merged model’s activations and the individual models’ activations. (5) Ties-Merging, which enhances merging by eliminating redundant parameters and resolving sign conflicts; (6) PCB-Merging, which balances parameter competition through intra-task significance and inter-task similarity analysis. We also report results from individual fine-tuned models and a pre-trained model on all tasks; (7) MetaGPT Zhou et al. (2024), which formalizes the objective of model merging into a multi-task learning framework, aiming to minimize the average loss difference between the merged model and each individual task model; (8) Knots Stoica et al. (2024), which utilizes Singular Value Decomposition to transform and align task-updates from multiple LoRA models into a shared representational space; (9) Consensus TA Wang et al. (2024) and (10) Localize-and-Stitch He et al. (2024), both of which focus on constructing a sparse mask for each task to reduce inter-task conflict during merging.
@@ 4.2 Results
Test Set Performance
Method Average paws qasc quartz story_cloze wiki_qa winogrande wsc
Pre-trained 53.5 49.9 35.8 53.3 48.1 76.2 50.0 61.1
Fine-tune 81.0 91.4 95.5 79.1 79.6 95.2 62.8 63.6
Parameter Average 60.2 55.2 57.5 55.2 49.4 91.1 50.2 62.5
Fisher Merging 68.2 67.9 84.4 63.5 57.1 90.1 54.2 60.8
RegMean 71.5 76.2 92.8 62.6 63.6 89.4 57.4 58.3
Task Arithmetic 71.6 71.1 81.4 62.1 77.1 95.0 57.4 56.9
Ties-Merging 72.2 78.8 88.4 65.1 70.7 84.2 56.2 62.3
PCB-Merging 72.3 71.5 91.7 66.8 62.7 92.8 57.1 63.3
MetaGPT 58.5 56.7 67.3 58.1 45.4 80.2 55.5 46.2
Consensus TA 72.3 73.8 83.4 61.9 76.0 83.8 56.2 61.0
Localize-and-Stitch 72.9 75.7 88.6 62.3 73.6 93.9 55.9 60.6
STF (ours) 73.7 77.4 89.1 62.6 75.2 94.2 56.4 61.1
Table 1: Merge fully-fine-tuned T5-base model with different methods.
Fully Fine-tuned NLP Model: T5 Merging
For NLP experiments, we evaluate on T5-base Raffel et al. (2020), an encoder-decoder transformer Vaswani et al. (2017) pre-trained with masked language modeling. We finetune T5-base on seven diverse tasks spanning question answering (QASC Khot et al. (2020), WikiQA Yang et al. (2015), QuaRTz Tafjord et al. (2019)), paraphrase identification (PAWS Zhang et al. (2019)), sentence completion (Story Cloze Sharma et al. (2018)), and coreference resolution (Winogrande Sakaguchi et al. (2020), WSC Levesque et al. (2012)). We use the code from Yadav et al. (2023) to finetune the model on these tasks. To reduce variance and ensure reliable evaluation, we report the experimental results of the average performance over different templates Bach et al. (2022) for each task.
As shown in Table˜1, STF achieves state-of-the-art performance when merging T5-base models, outperforming existing methods by $1.4\%$ on average across 7 tasks. STF shows particularly strong performance on PAWS and Story Cloze, with improvements of $5.9\%$ and $12.5\%$ respectively over PCB-Merging, the previous best method.
PEFT of NLP Model: LORA Adapters Merging
For PEFT experiments, we evaluate GPT-2 Medium model on LoRA adapters Hu et al. (2021), which are task-specific adapters that are fine-tuned on NLP task datasets. These tasks include converting tables (E2ENovikova et al. (2017)), knowledge graph (WebNLGGardent et al. (2017)) and structured data (DARTNan et al. (2020)) to natural language. We use the released checkpoints from Hu et al. (2021) and evaluate the performance of the merged model on these datasets. Specifically, we compare Knots, which is designed specifically for LoRA merging.
As shown in Table˜2, STF outperforms existing baselines on all metrics. Specifically, comparing with previous best baselines, STF shows improvements of $2.2\%$ for NIST of E2E, $4.8\%$ over PCB-Merging for CIDEr of E2E.
Test Set Performance
E2E DART WebNLG
[…]
Pre-trained 0.2 0.58 1.5 5.2 0.002 0.2 2.0 152.4 0.15 2.0 179.1
LoRA 67.7 8.64 46.0 68.3 2.36 44.8 35.0 50.4 52.3 37.0 44.4
Parameter Average 63.4 8.00 40.8 66.6 2.01 40.0 32.0 53.7 43.9 32.0 49.2 7
Task Arithmetic 63.2 8.02 40.9 66.3 1.98 40.8 33.0 53.7 45.9 34.0 48.8 6
Ties-Merging 62.8 8.14 41.1 65.9 2.08 41.4 33.0 53.7 46.1 33.0 48.1 5
PCB-Merging 62.9 8.12 41.4 66.0 2.08 41.3 33.0 53.5 46.2 33.0 48.1 3
Knots 62.9 8.22 41.6 65.8 2.05 41.52 34 54.0 46.0 34 47.5 2
MetaGPT 63.4 7.88 40.7 66.1 2.01 42.23 33 52.1 45.9 33 47.4 4
STF (ours) 64.1 8.40 42.2 66.5 2.18 41.6 33.0 54.1 47.1 34.0 48.0 1
Datasets
Method CMMLU GSM8K Humaneval Average
Chinese 38.6 2.3 13.4 18.1
Math 31.2 65.6 0 32.3
Code 33.3 0 17.1 16.8
Parameter Average 35.6 48.5 6.7 30.3
Task Arithmetic 35.4 46.1 9.8 30.4
Ties-Merging 36.5 53.4 12.8 34.3
MetaGPT 36.2 50.6 16.9 34.6
PCB-Merging 36.4 52.3 16.5 35.1
STF (ours) 36.5 63.0 14.0 37.8
Table 3: Results on the CMMLU, GSM8K, and Humaneval datasets.
Large Model: LLM Merging
We evaluate model merging on three fine-tuned Llama-2-7B models Touvron et al. (2023) focusing on different capabilities: Chinese language proficiency [^1] https://huggingface.co/LinkSoul/Chinese-Llama-2-7b, mathematical reasoning Yu et al. (2023) [^2] https://huggingface.co/meta-math/MetaMath-7B-V1.0, and code generation Rozière et al. (2023) [^3] https://huggingface.co/qualis2006/llama-2-7b-int4-python-code-18k. Each model’s performance was assessed using domain-specific benchmarks: CMMLU Li et al. (2023) for Chinese language understanding, GSM8K Cobbe et al. (2021) for mathematical reasoning, and HumanEval Chen et al. (2021) for code generation abilities.
As shown in Table˜3, STF achieves state-of-the-art performance across all three domains, improving overall performance by $2.7\%$ compared to the best baseline. The most significant improvement is observed in mathematical reasoning, where STF outperforms other methods by approximately $10\%$ on the GSM8K benchmark.
Out-of-Distribution Generalization
To evaluate out-of-distribution generalization, we test the merged T5-base model (previously trained on seven tasks) on six held-out tasks from the T0 mixture Sanh et al. (2021): three question answering tasks (Cosmos QA Huang et al. (2019), Social IQA Sap et al. (2019), QuAIL Rogers et al. (2020)), word sense disambiguation (WiC Pilehvar and Camacho-Collados (2018)), and two sentence completion tasks (COPA Roemmele et al. (2011), H-SWAG Zellers et al. (2019)). As shown in Table˜5, STF achieves $0.9\%$ improvement over the best baseline on T5-base, demonstrating strong generalization capabilities to tasks outside the training distribution.
4.3 Additional Results and Analysis
(a) Merge task matrices vs fine-tuned matrices
(b) Performance vs number of tasks and scaling factor
(c) Various $\eta$ at $\gamma$ = 0.6
(d) Various $\gamma$ at $\eta$ = 0.2
Figure 3: Experimental results on analyzing STF.
Task Matrices versus Fine-tuned Matrices
We compare merging task matrices versus fine-tuned matrices directly on T5-base models. As shown in Fig.˜3(a), merging task matrices consistently outperforms merging fine-tuned matrices across all architectures by a large margin. This validates our discussion in Section˜3.5 that task matrices better isolate task-specific features compared to fine-tuned matrices, making them more effective for model merging.
Feature Interference and Scaling Factor
We examine how performance of T5 merging varies with different number of tasks and scaling factor in Fig.˜3(b). We observe that as more tasks are merged, the optimal scaling factor $\gamma$ that achieves the best performance decreases. This trend indicates that larger $\gamma$ values amplify interference between task-specific features when merging many tasks, requiring smaller scaling factors to maintain performance. The results validate our discussion in Section˜3.5 about scaling merged matrix to balance feature preservation against interference.
Hyperparameter Sensitivity
We analyze the sensitivity of STF to the hyperparameters $\eta$ and $\gamma$ on the T5-base. Three model merging methods are compared in the experiments: STF, Ties-Merging, and PCB-Merging. We follow the experiment setup of previous work Du et al. (2024) to analyze the effect of two parameters. The first parameter is the scaling factor $\gamma$ , which is used to scale the merged task vector. For STF and Ties-Merging, the second hyperparameter is the top-k percentage $\eta$, indicating that the $\eta\%$ parameters in task vector with the highest absolute value are retained. For PCB-Merging, the second hyperparameter is mask ratio $\eta$, indicating that the $\eta\%$ parameters in parameter competition balancing matrix with the biggest absolute value will be retained. We vary the hyperparameters $\gamma$ with step size 0.1 from 0.5 to 1.0, and $\eta$ with step size 0.05 from 0.15 to 0.5.
First, in Figs.˜3(c) and 3(d), STF achieves optimal performance with $\gamma=0.8$ and $\eta=0.2$ on T5-base. Second, performance varies with both hyperparameters. For sparsity ratio $\eta$, performance first improves then declines as it increases. When $\eta$ is too low, important features are filtered out as most parameters are set to zero. When $\eta$ is too high, noise in the task matrices affects the quality of singular vectors. For scaling factor $\gamma$, we observe a similar trend - performance initially improves then deteriorates with increasing values. A small $\gamma$ diminishes the magnitude of singular vectors, while a large $\gamma$ may amplify interference between tasks.
Model Merging Time
We analyze the time required for model merging with STF on different models. As shown in Table˜4, the merging time increases with model size, primarily due to the SVD computation on larger matrices. For smaller models, TIES and PCB are more efficient than Knots and STF. However, with larger models, STF becomes more efficient because TIES and PCB require storing all parameters in memory simultaneously and run only on CPU, while STF processes one task matrix at a time and leverages GPU acceleration. The relatively short merging times demonstrate that STF is practical for real-world applications, despite its theoretical time complexity.
Method T5-base GPT-2 LoRA Llama-2-7B
TIES 34 0.08 1280
PCB 113 0.09 1593
Knots 8
STF (Ours) 127 5 623
Table 4: Merging time for different models (unit: s). We run PCB and TIES on CPU because of the large memory requirement.
5 Conclusion
In this paper, we present STF, a novel model merging approach that preserves task-specific features in linear transformations through feature superposition. Extensive experiments demonstrate that STF consistently outperforms existing methods across different architectures and tasks. The success of STF demonstrates the value of leveraging working mechanisms of deep neural networks in model merging, rather than treating them at the parameter level.
Limitations
While STF effectively preserves features during merging, it does not explicitly identify task-specific features corresponding to semantic meanings. Future work could leverage mechanistic interpretability techniques like superposition analysis and sparse autoencoders to better isolate and preserve task-specific features in linear transformations. Additionally, STF currently relies on a simple scaling approach to manage interference between tasks. More sophisticated methods for analyzing feature importance and selectively removing interference from less critical features could further improve performance. These advances would help develop a more principled approach to preserving task capabilities during model merging. Another major limitation is the time complexity of SVD and solving the linear system in STF, which can be improved by using more efficient method for feature extraction in the future.
Acknowledgements
This work is supported by National Key Research and Development Program of China (Grant No.2023YFB2903904), National Natural Science Foundation of China (Grant No. 92270106), Beijing Natural Science Foundation (Grant No. 4242039), and CCF-Huawei Populus Grove Fund.
Appendix A Measure Feature Preservation
We measure feature preservation by $\left|\langle\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)},\bar{\mathbf{M}}\mathbf{v}_{i}^{(k)}-\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}\rangle\right|$, where $\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}$ and $\bar{\mathbf{M}}\mathbf{v}_{i}^{(k)}$ and $\mathbf{v}_{i}^{(k)}$ are the task-specific features identified in Section˜3.2, and $\bar{\mathbf{M}}$ is the merged task matrix for various merging methods without scaling. This measurement is based on the superposition objective in (1). Higher values indicate that the output of merged task matrix $\bar{\mathbf{M}}$ is less similar to the output feature vectors of original task matrix $\mathbf{M}_{i}$, and thus the feature preservation is worse.
We calculate the preservation for every feature vector in the original fine-tuned models across all datasets, and then average all these individual preservation values. In Fig.˜1(a), we choose the input linear transformation matrices of all FFN layers in T5-base model. We also measured the feature preservation for other layers, including the attention layers. Fig.˜4 shows the feature preservation for the output layers of FFN in T5-base model, which shows the same trend with Fig.˜1(a). We observe that STF exactly preserves the magnitudes and directions of output feature vectors, while Ties-Merging and PCB-Merging have much lower feature preservation. The results indicate that better feature preservation strongly correlates with better performance and validate our hypothesis that feature superposition is effective for model merging.
Figure 4: Feature preservation for FFN output layer in T5-base model
Appendix B Singular Value Decomposition of Task Matrices
Applying SVD to task matrix, we have $\mathbf{M}_{i}=\sum_{k=1}^{r_{i}}\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}\mathbf{v}_{i}^{(k)\top}$, where $\sigma_{i}^{(k)}$ is the $k$-th singular value, and $\mathbf{u}_{i}^{(k)}$ and $\mathbf{v}_{i}^{(k)}$ are the $k$-th left and right singular vectors of $\mathbf{M}_{i}$, respectively. These singular vectors serve as orthogonal basis for output $\mathbf{M}_{i}\mathbf{x}$ and input $\mathbf{x}$. Specifically, any input representation $\mathbf{x}\in\mathbb{R}^{n}$ can be decomposed as $\mathbf{x}=\sum_{k=1}^{r_{i}}w_{i,j}^{(k)}\mathbf{v}_{i}^{(k)}+\mathbf{b}$, where $\mathbf{b}$ is orthogonal to all right singular vectors. When $\mathbf{M}_{i}$ transforms $\mathbf{x}$, we get $\mathbf{M}_{i}\mathbf{x}=\sum_{k=1}^{r_{i}}\sigma_{i}^{(k)}w_{i,j}^{(k)}\mathbf{u}_{i}^{(k)}$ since $\mathbf{b}$ lies in the null space of $\mathbf{M}_{i}$. This shows that the input $\mathbf{x}$ and output $\mathbf{M}_{i}\mathbf{x}$ can be represented as a weighted sum of right and left singular vectors, respectively.
Appendix C Proof of Theorem˜1
Here we use $\mathbf{X}=[\mathbf{x}_{i}^{(k)\top}\mathbf{x}_{i^{\prime}}^{(k^{\prime})}]_{i,k;i^{\prime},k^{\prime}}$ to denote the matrix formed by enumerating over $i,k$ for the row index and $i^{\prime},k^{\prime}$ for the column index.
Proof.
The objective of feature superposition can be converted to
$\left<\mathbf{u}_{i}^{(k)},\mathbf{M}\mathbf{v}_{i}^{(k)}\right>=\left<\mathbf{u}_{i}^{(k)},\mathbf{M}_{i}\mathbf{v}_{i}^{(k)}\right>,\quad\forall i,k.$
We bring $\mathbf{M}=\sum\nolimits_{i^{\prime}=1}^{T}\sum\nolimits_{k^{\prime}=1}^{r_{i}}\alpha_{i^{\prime}}^{(k^{\prime})}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}$ and $\mathbf{M}_{i}=\sum\nolimits_{k^{\prime}=1}^{r_{i}}\alpha_{i}^{(k^{\prime})}\mathbf{u}_{i}^{(k^{\prime})}\mathbf{v}_{i}^{(k^{\prime})\top}$ to the left side of the equation, and move $\sigma_{i}^{(k)}$ to the right side.
Then, the objective becomes
$\mathbf{u}_{i}^{(k)\top}\!\!\left(\sum_{i^{\prime}=1}^{T}\sum_{k^{\prime}=1}^{r_{i}}\!\alpha_{i^{\prime}}^{(k^{\prime})}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\!\!\right)\!\mathbf{v}_{i}^{(k)}\!=\!\sigma_{i}^{(k)}\Leftrightarrow\sum_{i^{\prime}=1}^{T}\sum_{k^{\prime}=1}^{r_{i}}\!\!\mathbf{u}_{i}^{(k)\top}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\mathbf{v}_{i}^{(k)}\alpha_{i^{\prime}}^{(k^{\prime})}\!=\!\sigma_{i}^{(k)}$ (4)
which is one linear equation with $\mathbf{u}_{i}^{(k)\top}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\mathbf{v}_{i}^{(k)}$ as the coefficient of variable $\alpha_{i^{\prime}}^{(k^{\prime})}$. We can convert all equations in (1) to a linear system for all tasks in $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$:
$\begin{bmatrix}\mathbf{u}_{1}^{(1)\top}\mathbf{u}_{1}^{(1)}\mathbf{v}_{1}^{(1)\top}\mathbf{v}_{1}^{(1)}&\cdots&\mathbf{u}_{1}^{(1)\top}\mathbf{u}_{T}^{(r_{T})}\mathbf{v}_{T}^{(r_{T})\top}\mathbf{v}_{1}^{(1)}\\ \vdots&\ddots&\vdots\\ \mathbf{u}_{T}^{(r_{T})\top}\mathbf{u}_{1}^{(1)}\mathbf{v}_{1}^{(1)\top}\mathbf{v}_{T}^{(r_{T})}&\cdots&\mathbf{u}_{T}^{(r_{T})\top}\mathbf{u}_{T}^{(r_{T})}\mathbf{v}_{T}^{(r_{T})\top}\mathbf{v}_{T}^{(r_{T})}\end{bmatrix}\begin{bmatrix}\alpha_{1}^{(1)}\\ \vdots\\ \alpha_{T}^{(r_{T})}\end{bmatrix}=\begin{bmatrix}\sigma_{1}^{(1)}\\ \vdots\\ \sigma_{T}^{(r_{T})}\end{bmatrix}$.
With Kronecker product, we can write the linear system in a more compact form:
$\mathbf{\Sigma}\circ\mathbf{U}\circ\mathbf{V}\boldsymbol{\alpha}=\boldsymbol{\sigma},$
where $\mathbf{U}=[\mathbf{u}_{i}^{(k)\top}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}]_{i,k;i^{\prime},k^{\prime}}=\bar{\mathbf{U}}^{\top}\bar{\mathbf{U}}\in\mathbb{R}^{r\times r}$; $\mathbf{V}=[\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\mathbf{v}_{i}^{(k)}]_{i^{\prime},k^{\prime};i,k}=\bar{\mathbf{V}}^{\top}\bar{\mathbf{V}}\in\mathbb{R}^{r\times r}$; $\boldsymbol{\alpha}$ is the vector of variables $\alpha_{i^{\prime}}^{(k^{\prime})}$; and $\boldsymbol{\sigma}$ is the vector of singular values $\sigma_{i}^{(k)}$. The index $i,k,i^{\prime},k^{\prime}$ share consistent order across $\mathbf{U}$ and $\mathbf{V}$. ∎
Appendix D More Experimental Results
Out-of-Distribution Performance
Method Average cosmos_qa social_iqa quail wic copa h-swag
Pre-trained 31.1 21.9 18.8 24.1 65.6 43.8 12.5
Average 36.3 23.5 37.0 25.4 50.2 54.5 27.2
Task Arithmetic 37.0 21.9 36.8 25.5 49.5 61.4 26.6
Ties-Merging 37.1 22.2 37.8 24.9 51.6 62.2 26.5
PCB-Merging 37.3 23.0 38.1 24.6 52.2 59.0 27.2
STF (ours) 38.2 22.4 38.0 26.0 51.6 64.1 27.0
D.1 More Analysis
Removing Singular Vectors
We investigate removing singular vectors from task matrices before merging, analyzing on T5-base: removing smallest versus largest singular vectors. As shown in Figs.˜5(b) and 5(a), removing up to 80% of smallest singular vectors has minimal impact on performance, while removing just 5% of largest singular vectors causes significant degradation. This indicates that while most singular features are redundant, preserving those with largest singular values is crucial for maintaining model capabilities. This insight could guide future work on selective feature preservation during model merging.
(a) Remove smallest singular vectors
(b) Remove biggest singular vectors
Figure 5: Performance of STF with different percentage of removed features.
Fully Fine-tuned Vision Model: ViT Merging
We evaluate STF on vision tasks using two CLIP models Radford et al. (2021) with ViT-B/32 and ViT-L/14 architectures Dosovitskiy (2020) as visual encoders. We use the released checkpoints from Ilharco et al. (2022) that were fine-tuned on eight diverse classification tasks spanning remote sensing, traffic signs, and satellite imagery domains. These tasks include: Cars Krause et al. (2013), DTD Cimpoi et al. (2014), EuroSAT Helber et al. (2019), GTSRB Stallkamp et al. (2011), MNIST LeCun (1998), RESISC45 Cheng et al. (2017), SUN397 Xiao et al. (2016), and SVHN Netzer et al. (2011). For all experiments, we keep the text encoder fixed and only merge the parameters of visual encoders.
As shown in Table˜6, STF achieves state-of-the-art performance when merging fully fine-tuned ViT models, outperforming existing methods by 1.1% and 0.3% on average across 8 tasks for ViT-B/32 and ViT-L/14 architectures respectively. For ViT-B/32, STF shows particularly strong performance on RESISC45 and EuroSAT, with 2.9% and 11.1% improvements over the best baseline. For ViT-L/14, STF maintains high performance across all tasks, demonstrating effective preservation of task-specific features during merging.